# Collective Noise Filtering in Complex Networks

### v1

@@ Abstract
Complex networks are powerful representations of complex systems across scales and domains, and the field is experiencing unprecedented growth in data availability. However, real-world network data often suffer from noise, biases, and missing data in the edge weights, which undermine the reliability of downstream network analyses. Standard noise filtering approaches, whether treating individual edges one-by-one or assuming a uniform global noise level, are suboptimal, because in reality both signal and noise can be heterogeneous and correlated across multiple edges. As a solution, we introduce the Network Wiener Filter, a principled method for collective edge-level noise filtering that leverages both network topology and noise characteristics, to reduce error in the observed edge weights and to infer missing edge weights. We demonstrate the broad practical efficacy of the Network Wiener Filter in two distinct settings, the genetic interaction network of the yeast S. cerevisiae and the Enron Corpus email network, noting compelling evidence of successful noise suppression in both applications. With the Network Wiener Filter, we advocate for a shift toward error-aware network science, one that embraces data imperfection as an inherent feature and learns to navigate it effectively.
@@
Ever since the early development of graph theory, mathematicians have expanded its toolkit far beyond elegant puzzles, building increasingly sophisticated theories for reasoning about connectivity and structure [14]. In recent decades, network science transformed these abstract ideas into the empirical world, using networks to represent complex systems across diverse disciplines, including, among others, social science [36], biology [3], transportation [5], climate science [21], finance and economics [1]. Analyzing these complex systems from a network perspective has enabled researchers to uncover hidden patterns, capture emergent properties, and generate unique insights [4]. While the network science community has benefited from a rapid expansion in the availability of network data, it is frequently overlooked that such data can be imperfect. Real-world networks are often subjected to various kinds of errors such as noise, biases, and incomplete observation [41, 33, 58, 19]. Yet it remains common in the literature to omit any quantitative assessment or reporting of these errors [46]. Moreover, most network analysis pipelines do not consider potential data errors and treat empirical networks as if they were exact graphs, in the spirit of graph theory. Ignoring the issue of data reliability in this manner can compromise the robustness of downstream network-based inference, leading to flawed conclusions and decisions.
@@
In response, a growing body of work has sought to mitigate edge noise in network data by exploiting network topological information, including link prediction methods [37, 18, 64, 38, 7, 39, 63, 32, 60], diffusion-based methods [22, 59, 62], generative model-based methods [15, 28, 45, 44, 50, 49], and singular value shrinkage methods [16, 9, 52, 43, 24, 29, 25, 55]. Each method category relies on different, often strong, network structural assumptions and is tailored to different use cases (SI Table [placeholder]). However, none of these approaches is adequate for the realistic scenario of having heterogeneous and correlated noise. Either noise level is not explicitly incorporated, as in link prediction and diffusion-based methods, or noise is assumed to be largely homogeneous and independent across edges, as in generative model-based and singular value shrinkage methods. In the case of heterogeneous, or even correlated, noise, practitioners often resort to local thresholding strategies at the level of individual edges, without harnessing the rich information in the network architecture [19]. This, in turn, can limit the amount of signal that can be uncovered from the noisy data.
@@
In this work, we propose the Network Wiener Filter (NetWF) as a principled method for collective noise filtering in complex networks. Grounded in the classical theory of Wiener filtering, the NetWF takes into account network topology through an intuitive edge-similarity measure, which is then combined with explicit edge-level noise statistics in a synergistic way. Importantly, this formulation naturally accommodates heterogeneous or correlated noise across edges, as well as the special case of homogeneous and independent noise, rendering it fully general with respect to noise structure. At the same time, the NetWF applies seamlessly to directed or undirected, binary, weighted or even signed networks. As illustrated on our real-world applications, our efficient implementation using the conjugate gradient algorithm enables the NetWF to scale to large empirical networks.
@@
In classical signal processing, the Wiener filter is a foundational framework designed to recover the underlying true signal from noisy observations by minimizing the expected mean square error (MSE) between the estimate and the (unknown) true signal [47]. Using the statistical properties of both the signal and the noise, it seeks an optimal balance between noise suppression and signal preservation. As a result, Wiener filtering has been widely adopted for denoising tasks such as in time series (Figure 1A) and image processing (Figure 1B).
@@
Here, we start with a generalized Wiener filter formalism [51] and extend it to network data. Imagine that we observe a demeaned data vector $a=u+n\in\mathbb{R}^{k}$, where $u$ denotes the true signal and $n$ the additive noise. For data types that are not inherently one-dimensional, $a$ can be understood as a vectorized representation, such as a vectorized image or a vectorized network adjacency matrix, $a=\mathrm{vec}(A)$, where $A_{ij}$ stands for the interaction weight between nodes $i$ and $j$ in the network. The goal is then to identify a “Wiener operator” $G^{\mathrm{W}}$ that, when acting on $a$, gives an estimate of the signal $\hat{u}=G^{\mathrm{W}}a$ that minimizes the mean square error (MSE):
@@
where the expectation $\mathbb{E}$ is taken over the joint realizations of $u$ and $n$, and $\lVert\cdot\rVert_{2}$ is the $L^{2}$ norm. The optimization problem of Eq. 1 admits a closed-form analytical solution, see SI Text [placeholder]. In the case when $u$ and $n$ are drawn from independent distributions, the solution is given by
@@
in terms of the covariance matrices $C_{u}=\mathbb{E}[uu^{T}]$ and $C_{n}=\mathbb{E}[nn^{T}]$, both with dimensionality $\mathbb{R}^{k\times k}$.
@@
Figure 1: Illustration of Wiener filtering for noise reduction across data types, and why networks are different. Starting from (A) a time series, (B) an image, and (C) a bipartite toy network, we add independent and identically distributed Gaussian white noise to obtain noisy observations. We then apply Wiener filtering to estimate the underlying signal from noisy observations using the known noise statistics (more details in SI Text [placeholder] and SI Figure [placeholder]). In time series and images, the classical Wiener filter exploits data-independent temporal or spatial proximity to capture signal correlations, with the intuition that nearby samples tend to be similar. In contrast, network data is inherently high-dimensional and requires a data-dependent adaptive treatment. The NetWF bridges this gap by using a network-informed edge-similarity measure that captures correlations globally across edges, as discussed in the main text. In the toy network, our NetWF reveals the original structure of two communities.
@@
Applying the Wiener filter hinges on having access to the two covariance matrices, or, more realistically, reliable proxies for them. $C_{n}$ reflects noise characteristics and typically must be supplied through prior knowledge of how noise enters the data. In the simplest case, noise is independent and homogeneous, and $C_{n}$ reduces to a scaled identity matrix, as we imposed in our illustrative examples in Figure 1. More generally, noise can be heterogeneous across entries, or even correlated, in which case $C_{n}$ is a generic, positive semidefinite covariance matrix. On the other hand, $C_{u}$ boils down to signal variability and correlation, where the latter can be interpreted as the intrinsic similarity structure of the signal. For one-dimensional time series and two-dimensional images, this similarity structure is guided by proximity, as nearby samples are more similar than distant ones. Note that all images share the same grid-like proximity structure, independent of the actual pixel values. Then it is straightforward to specify a similarity structure that respects this regular grid architecture of the data, often a priori by positing a stationary kernel. It is also convenient to work in the frequency domain via the Fourier transform, where $C_{u}$ can be described compactly by a power spectral density through the Wiener–Khinchin theorem, achieving high computational efficiency for the Wiener filter [47].
@@
In contrast, networks (Figure 1C) are fundamentally different: the data is inherently high-dimensional with intricate correlations that depend on the input data. To provide the required context for Wiener filtering, we must construct an edge-similarity measure that serves as the proxy for signal correlation in a qualitatively different, data-adaptive way, one that exploits the observed network patterns. In the proposed implementation of the NetWF, we do so using the intuition that edges between similar pairs of nodes are expected to be similar, as discussed next. As a result, in the example of Figure 1C the NetWF successfully reveals the original structure of two network communities.
@@
For the purpose of community detection, a local notion of edge-similarity has been already proposed for edges that share an endpoint [2]. The fundamental idea is that any notion of node similarity can be used between the non-shared endpoints of the two edges to quantify the similarity of the edge pair. Here, we go one step beyond this local notion of edge-similarity and introduce a global edge-similarity between any two edges in the network, starting with the case of directed networks (Figure 2A). For two directed edges $\overrightarrow{AB}$ and $\overrightarrow{CD}$, we separately model the similarity between the two source nodes, $A$ and $C$, and between the two target nodes, $B$ and $D$. As a convenient choice that works for any (binary, weighted or signed) networks, we quantify these node similarities using the Pearson correlation between their connection profiles. Specifically, the source profile similarity compares outgoing connection patterns,
@@
Although the Wiener filter appears to be a linear framework with respect to input data, the filtering process is actually nonlinear if the Wiener operator $G^{\mathrm{W}}$ of Eq. 2 itself is data-dependent, as is the case for the proposed NetWF. For detailed computational implementation of the NetWF, including a conjugate gradient-based iterative implementation that makes handling large empirical networks feasible, see the Materials and Methods.
@@
Figure 3: Evidence that the NetWF reduces experimental noise in the yeast S. cerevisiae genetic interaction (GI) network. (A) Validation of negative GI values against three biological benchmarks: protein–protein interaction (PPI), protein co-complex membership, and Gene Ontology (GO) co-annotation. Relative to the raw data and the optimal shrinker (OS) baseline, NetWF improves both area under the precision–recall curve (AUPRC) and fold enrichment across all benchmarks. (B) A 10-fold cross validation (CV) test: in each fold, a random subset of GI entries is masked and predicted from the remaining data. The NetWF and the OS achieve significantly lower average mean square error (MSE) than naïve mean imputation (MI), where the missing entries are imputed using the mean of available GI values in each gene. Error bars indicate standard errors across folds. (C) Example of inferring a novel GI within the exocyst complex (SEC15–SEC6) by the NetWF. Dashed edges highlight a few example profile similarities within the complex that facilitate information propagation, collectively shifting the GI value from zero to a strongly negative estimate. (D) Profile similarity networks (PSNs) showcasing similarities between gene pairs (Eq. 6), before and after the NetWF is applied to the raw GI network. Edge width scales with similarity, and similarities below $0.2$ are omitted for clarity.
@@
Large-scale biological data are inherently noisy and incomplete, yet we rarely have detailed, quantitative insights into how much error there is in the measurements. A rare exception is the genome-wide pairwise genetic interaction (GI) network of the baker’s yeast S. cerevisiae [19]. Pairwise GIs arise when the fitness phenotype of a double mutant $f_{i,j}$ deviates from the expectation corresponding to independent single mutants fitness values, $f_{i}$ and $f_{j}$:
@@
Here, the fitness is measured as colony growth relative to a wild-type (non-mutant) reference under standardized conditions, via high-throughput synthetic genetic array (SGA) analysis [6]. Notably, as a result of the experiments, not only the $\epsilon_{i,j}$ edge weights are provided for each gene pair, but also the experimental error bar for each weight [19]. Most of the interaction weights were found to be comparable to the noise level, indicating that the independent baseline fitness expectation is valid for most gene pairs. A negative GI ($\epsilon_{i,j}<0$) indicates that the double mutant grows worse than expected, in extreme cases known as synthetic lethality. Conversely, a positive GI ($\epsilon_{i,j}>0$) indicates that the double mutant grows better than expected, i.e. genetic buffering or suppression. To identify true GIs, the authors suggested three alternative filtering approaches (lenient, intermediate, stringent) to the edge level data based on a combination of effects size ($\epsilon_{i,j}$) and noise reporting ($p$-value).
@@
These systematic fitness screens uncovered a rich hierarchical functional organization in the eukaryotic cell, covering most of the pairwise GI space between $\sim 6{,}000$ genes in S. cerevisiae. The organization of GIs was found to be qualitatively different for essential genes (required for growth) vs. non-essential genes, as well as for positive vs. negative GIs. Here, as an illustration of the NetWF, we restrict our analysis to the dense part of the network between essential genes (ExE), comprising only temperature-sensitive alleles (427 nodes), see Materials and Methods for details. Besides providing rich functional information, it was also found that GIs are enriched in additional biological information. In the ExE space, this is manifested in the significant overlap between negative GIs and other data sources, including protein–protein interactions (PPI) [48], protein co-complex membership [6, 10], and Gene Ontology (GO) co-annotation [17], see Materials and Methods for details. The relatively few cases of extreme positive GIs also provide information on genetic function; however, positive GIs were deemed to be somewhat problematic, potentially being diluted with false positives, prompting focused experimental efforts on genetic suppression [56, 57].
@@
Using a stringent weight threshold, $\epsilon<-0.12$ (negative) and $\epsilon>0.16$ (positive), the filtered ExE map contains $274$ positive GIs and 6,7056{,}705 negative GIs. The noise variance information entails heterogeneous but uncorrelated edge-level noise, corresponding to a diagonal noise covariance matrix $C_{n}$. After the NetWF denoising (SI Figure [placeholder]), using the stringent weight thresholds yields only $21$ positive GIs ($96\%$ reduction in count) and 4,8384{,}838 negative GIs ($28\%$ reduction in count), including $716$ new negative GIs compared to the local filtering. Notably, the NetWF reduces disproportionally the number of (potentially problematic) positive GIs, while even unveiling previously neglected negative GIs. More importantly, the filtered ExE network demonstrates stronger adherence to all three benchmarks (PPI, co-complex, GO), improving both area under the precision–recall curve (AUPRC) and fold enrichment relative to the raw data (Figure 3A). Here, fold enrichment compares the representation of benchmark pairs among the 1,0001{,}000 most negative GIs to their background frequency across all ExE pairs.
@@
Next, as an illustration of link prediction performance, we carry out a 10-fold cross validation test (Figure 3B). We divide the observed symmetric GI entries into 10 partitions, and in each fold, we mask the partitioned entries to missing values and predict them from the remaining observations, see Materials and Methods for details. The NetWF delivers substantially lower MSE than a naïve mean imputation that fills missing entries using the mean of available GI values for each gene.
@@
As a case study, we successfully recover a missing interaction from the input data within the exocyst complex [42]. The NetWF infers a strong negative GI between genes SEC15–SEC6 ($\epsilon=-0.23$), in agreement with the known biophysical interaction [23, 54]. Although the prediction originates from the NetWF processing collective information from the entire network, we provide an interpretation within the exocyst subgraph (Figure 3C). Notably, SEC15 is highly similar to SEC10 and SEC8 in GI profile, while SEC6 is similar to SEC5; therefore the strong negative GI of SEC10–SEC5 and SEC8–SEC5 contributes to the negative SEC15–SEC6 estimate. See SI Figure [placeholder] for the physical structure of the exocyst complex.
@@
Finally, to quantify gene-level similarity, we construct profile similarity networks (PSNs) in which edge weights are Pearson correlations between GI profiles (Eq. 6). Using the originally proposed cutoff >0.2>0.2 [19], the PSN has 3,6863{,}686 edges before filtering and 11,10811{,}108 edges after filtering (Figure 3D). This three-fold increase indicates fewer idiosyncratic differences in gene connection profiles, i.e. a smaller effective degrees of freedom. Together with the evidence that biological signal is preserved or even enhanced (Figure 3A), this supports a higher signal-to-noise ratio after filtering.
@@
In comparison, the OS baseline behaves inconsistently across evaluations. It adheres worse to the PPI and protein co-complex benchmarks than even the raw data (Figure 3A), yet performs well in cross validation (Figure 3B), yields a dense PSN (18,24718{,}247 edges), and also recovers the SEC15–SEC6 interaction ($\epsilon=-0.17$).
@@
To test how much the performance of the NetWF depends on having access to the rich heterogeneous noise information, we run a restricted, homogenized-noise version in which we replace the heterogeneous variances by their mean, just like for the OS. The restricted results across the board change only marginally compared to those with the full NetWF, with over a $99\%$ overlap in terms of negative GIs (SI Figure [placeholder]). This indicates that the NetWF can be a useful tool across various applications, even without detailed noise characterization, as also illustrated in our next application.
Figure 4: Evidence that the NetWF reduces sampling noise in the Enron Corpus email network. (A) Full-year email exchange frequency network of 2001 between employees as nodes. Directed edges point from sender to recipient, with edge weight equal to email frequency (defined as the number of emails per month). Node size reflects the sum of incoming and outgoing weights. For clarity, edges with weight <0.2<0.2 are not shown. (B) Raw October 2001 snapshot, treated as a noisy calendar-month sample of the full-year network, exhibiting pronounced fluctuations in edge weights. (C) October 2001 after applying the NetWF, with the extreme weights reduced without compromising the global structure. (D) Relative to the raw October network and the OS baseline, the NetWF largely reduces the October MSE with respect to the full-year network. Repeating for all twelve calendar-month snapshots, the NetWF also attains the lowest and most stable average MSE. Error bars indicate standard errors across calendar months.
@@
The Enron Corpus is a publicly available database containing internal email exchanges among key Enron employees in the years leading up to the company’s famous bankruptcy in December 2001, and it has become a widely used resource for studying organizational communication patterns [31]. We use this dataset for two complementary goals. First, we demonstrate the performance of the NetWF in the case of heterogeneous and correlated noise covariance $C_{n}$, the highest level of noise characteristics. Second, we take the opportunity to embrace the more common empirical setting where no noise statistics is provided, and stress test the extent to which the NetWF can perform denoising under this far from optimal condition.

### v2

@@ Abstract
Complex networks are powerful representations of complex systems across scales and domains, and the field is experiencing unprecedented growth in data availability. However, real-world network data often suffer from noise, biases, and missing data in edge weights, which undermine the reliability of downstream network analyses. Standard noise filtering approaches, whether treating individual edges one-by-one or assuming a uniform global noise level, are suboptimal, because in reality both signal and noise can be heterogeneous and correlated across multiple edges. As a solution, we introduce the Network Wiener Filter, a principled method for collective edge-level noise filtering that leverages both network structure and noise characteristics, to reduce error in the observed edge weights and to infer missing edge weights. We demonstrate the broad practical efficacy of the Network Wiener Filter in two distinct settings, the genetic interaction network of the budding yeast S. cerevisiae and the Enron Corpus email network, noting compelling evidence of successful noise suppression in both applications. With the Network Wiener Filter, we advocate for a shift toward error-aware network science, one that embraces data imperfection as an inherent feature and learns to navigate it effectively.
@@
Ever since the early development of graph theory, mathematicians have expanded its toolkit far beyond elegant puzzles, building increasingly sophisticated theories for reasoning about connectivity and structure [14]. In recent decades, network science has carried these abstract ideas into the empirical world, using networks to represent complex systems across diverse disciplines, including, among others, social science [36], biology [3], transportation [5], climate science [21], finance and economics [1]. Analyzing these complex systems from a network perspective has enabled researchers to uncover hidden patterns, capture emergent properties, and generate unique insights [4]. While the network science community has benefited from a rapid expansion in the availability of network data, it is frequently overlooked that such data can be imperfect. Real-world networks are often subjected to various kinds of errors such as noise, biases, and incomplete observation [41, 33, 58, 19]. Yet it remains common in the literature to omit any quantitative assessment or reporting of these errors [46]. Moreover, most network analysis pipelines do not consider potential data errors and treat empirical networks as if they were exact graphs, in the spirit of graph theory. Ignoring the issue of data reliability in this manner can compromise the robustness of downstream network-based inference, leading to flawed conclusions and decisions.
@@
In response, a growing body of work has been seeking to mitigate edge noise in network data by exploiting network topological information, including link prediction methods [37, 18, 64, 38, 7, 39, 63, 32, 60], diffusion-based methods [22, 59, 62], generative model-based methods [15, 28, 45, 44, 50, 49], and singular value shrinkage methods [16, 9, 52, 43, 24, 29, 25, 55]. Each method category relies on distinct, often strong, network structural assumptions and is tailored to different use cases (SI Table I). However, none of these approaches is adequate for the realistic scenario of having heterogeneous and correlated noise. Either noise level is not explicitly incorporated, as in link prediction and diffusion-based methods, or noise is assumed to be largely homogeneous and independent across edges, as in generative model-based and singular value shrinkage methods. In the case of heterogeneous, or even correlated, noise, practitioners often resort to local thresholding strategies at the level of individual edges, without harnessing the rich information in the network architecture [19]. This, in turn, can limit how effectively true signal can be extracted from the noisy data.
@@
In this work, we propose the Network Wiener Filter (NetWF) as a principled method for collective noise filtering in complex networks. Grounded in the classical theory of Wiener filtering [61], the NetWF takes into account the network structure through an intuitive edge-similarity measure, which is then combined with explicit edge-level noise statistics in a synergistic way. Importantly, this formulation naturally accommodates heterogeneous noise or noise correlated across edges, as well as the special case of homogeneous and independent noise, rendering it fully general in terms of noise structure. At the same time, the NetWF applies seamlessly to directed or undirected, binary, weighted or even signed networks. As demonstrated with our real-world applications, our efficient implementation using the conjugate gradient algorithm enables the NetWF to scale to large empirical networks.
@@
In classical signal processing, the Wiener filter is a foundational framework designed to recover the underlying true signal from noisy observations by minimizing the expected mean squared error (MSE) between the estimate and the (unknown) true signal [61, 47]. Using the statistical properties of both the signal and the noise, it seeks an optimal balance between noise suppression and signal preservation. As a result, Wiener filtering has been widely adopted for denoising tasks such as in time series (Figure 1A) and image processing (Figure 1B).
@@
Here, we start with a generalized Wiener filter formalism [51] and extend it to network data. Imagine that we observe a data vector $a=u+n\in\mathbb{R}^{k}$, where $u$ denotes the true signal and $n$ the additive noise; both of which are zero-mean random vectors drawn from unknown data generating processes. For data types that are not inherently one-dimensional, $a$ can be understood as a vectorized representation, such as a vectorized image or a vectorized network adjacency matrix, $a=\mathrm{vec}(A)$, where $A_{ij}$ stands for the interaction weight between nodes $i$ and $j$ in the network. The goal is then to identify a “Wiener operator” $G^{\mathrm{W}}$ that, when acting on $a$, gives an estimate of the signal $\hat{u}=G^{\mathrm{W}}a$ that minimizes the MSE with respect to the $u$:
@@
where the expectation $\mathbb{E}$ is taken over the joint realizations of $u$ and $n$, and $\lVert\cdot\rVert_{2}$ is the $L^{2}$ norm. Assuming known second-order statistics of the data generating processes governing $u$ and $n$, the optimization problem of Eq. 1 admits a closed-form analytical solution; see SI Text B for a full derivation. Specifically, in the case when $u$ and $n$ are drawn from independent distributions, the solution is given by
@@
in terms of the covariance matrices $C_{u}=\mathbb{E}[uu^{T}]$ and $C_{n}=\mathbb{E}[nn^{T}]$, both with dimensionality $\mathbb{R}^{k\times k}$, where the expectations are taken over the independent realizations of $u$ and $n$.
@@
Figure 1: Illustration of Wiener filtering for noise reduction across data types, and why networks are different. Starting from (A) a gravitational wave time series, (B) an image, and (C) a toy network with two communities, we add independent and identically distributed Gaussian white noise to obtain noisy observations. We then apply Wiener filtering to estimate the underlying signal from noisy observations using the known noise statistics (more details in SI Text A). In time series and images, the classical Wiener filter exploits data-independent temporal or spatial proximity to capture signal correlations, with the intuition that nearby samples tend to be similar. In contrast, network data is inherently high-dimensional and lacks a built-in notion of proximity, requiring a new, data-adaptive treatment. The NetWF bridges this gap by using a network-informed edge-similarity measure that captures correlations globally across edges, as discussed in the main text. In the toy network, our NetWF reveals the original structure of two communities (more quantitative results in SI Figure 1).
@@
Applying the Wiener filter hinges on having access to the two covariance matrices, or, more realistically, reliable proxies for them. $C_{n}$ reflects noise characteristics and typically must be supplied through prior knowledge of how noise enters the data. In the simplest case, noise is independent and homogeneous, and $C_{n}$ reduces to a scaled identity matrix, as we imposed in our illustrative examples in Figure 1. More generally, noise can be heterogeneous across entries, or even correlated, in which case $C_{n}$ is a generic, positive semidefinite covariance matrix. On the other hand, $C_{u}$ boils down to signal variability and correlation, where the latter can be interpreted as the intrinsic similarity structure of the signal. For one-dimensional time series and two-dimensional images, this similarity structure is guided by proximity, as nearby samples are more similar than distant ones. Note that all images share the same grid-like data architecture, independent of the actual pixel values. Then it is straightforward to specify a similarity structure that respects this regular geometry, often a priori by positing a stationary kernel. It is also convenient to work in the frequency domain via the Fourier transform, where $C_{u}$ can be described compactly by a power spectral density through the Wiener–Khinchin theorem, achieving high computational efficiency for the Wiener filter [47].
@@
In contrast, networks (Figure 1C) are fundamentally different: they are inherently high-dimensional objects whose intricate signal correlations must be inferred from the network data itself. To provide the required context for Wiener filtering, we construct an edge-similarity measure that serves as a proxy for signal correlation in a qualitatively different, data-adaptive way, one that exploits the observed network patterns. In the proposed implementation of the NetWF, we do so using the intuition that edges between similar pairs of nodes are expected to be similar, as discussed next. In the example of Figure 1C, the NetWF successfully reveals the original structure of two network communities.
@@
For the purpose of community detection, a local notion of edge-similarity has already been proposed for edges that share an endpoint [2]. The fundamental idea is that a suitable notion of node similarity can be used between the non-shared endpoints of the two edges to quantify the similarity of the edge pair. Here, we go one step beyond this local notion of edge-similarity and introduce a global edge-similarity between any two edges in the network, starting with the case of directed networks (Figure 2A). For two directed edges $\overrightarrow{AB}$ and $\overrightarrow{CD}$, we separately model the similarity between the two source nodes, $A$ and $C$, and between the two target nodes, $B$ and $D$. As a convenient choice that works for any (binary, weighted or signed) networks, we quantify these node similarities using the Pearson correlation between their connection profiles. Specifically, the source profile similarity compares outgoing connection patterns,
@@
Although the Wiener filter appears to be a linear framework with respect to input data, the filtering process is actually nonlinear if the Wiener operator $G^{\mathrm{W}}$ of Eq. 2 itself is data-dependent, as is the case for the proposed NetWF. For detailed computational implementation of the NetWF, including a conjugate gradient-based iterative implementation that makes handling large empirical networks feasible, see Materials and Methods.
@@
Large-scale biological data are inherently noisy and often incomplete, yet we rarely have detailed, quantitative insights into how much error there is in the measurements. An exception is the genome-wide pairwise genetic interaction (GI) network of the budding yeast S. cerevisiae [19]. Pairwise GIs arise when the fitness phenotype of a double mutant, $f_{i,j}$, deviates from the expectation under independent single-mutant effects, $f_{i}$ and $f_{j}$:
@@
Here, the fitness is measured as colony growth relative to a wild-type (non-mutant) reference under standardized conditions, via high-throughput synthetic genetic array (SGA) analysis [6]. Notably, the experiments not only provide the $\epsilon_{i,j}$ edge weights for each gene pair, but also the experimental error bar for each weight [19]. Most interaction weights were found to be comparable to the noise level, suggesting that the independent fitness expectation is valid for most gene pairs. A negative GI ($\epsilon_{i,j}<0$) indicates that the double mutant grows worse than expected, in extreme cases known as synthetic lethality. Conversely, a positive GI ($\epsilon_{i,j}>0$) indicates that the double mutant grows better than expected, i.e. genetic buffering or suppression. To identify true GIs, the authors proposed three thresholding schemes (lenient, intermediate, stringent) for edge-level data based on a combination of effect size ($\epsilon_{i,j}$) and noise reporting ($p$-value).
@@
These systematic fitness screens uncovered a rich hierarchical functional organization in an eukaryotic cell, covering most of the pairwise GI space between $\sim 6{,}000$ genes in S. cerevisiae. The organization of GIs was found to be qualitatively different for essential genes (required for growth) vs. non-essential genes, as well as for positive vs. negative GIs. Here, we focus our analysis on the dense part of the network between essential genes (ExE), comprising $855$ nodes; see Materials and Methods for details. Besides providing rich functional information, it was also found that GIs encode additional biological information [19]. In the ExE space, this is manifested in the significant overlap between negative GIs and other data sources, including protein–protein interactions (PPI) [48], protein co-complex membership [6, 10], and Gene Ontology (GO) bioprocess term co-annotation [17]; see Materials and Methods for details. The relatively few cases of extreme positive GIs also provide information on genetic function; however, positive GIs were deemed to be somewhat problematic, potentially being diluted with false positives, prompting additional focused experimental efforts on genetic suppression [56, 57].
@@
Using a stringent weight threshold [19], $\epsilon<-0.12$ (negative) and $\epsilon>0.16$ (positive), the ExE map contains 3,5163{,}516 positive GIs and 26,26826{,}268 negative GIs. The noise variance information entails uncorrelated but heterogeneous edge-level noise, corresponding to a diagonal noise covariance matrix $C_{n}$. In contrast, applying the NetWF denoising (SI Figure 2) and then applying the same stringent weight threshold yields only $620$ positive GIs ($82\%$ reduction in count) and 17,55817{,}558 negative GIs ($33\%$ reduction in count), including $2034$ new negative GIs. In other words, the NetWF disproportionately reduces the number of (potentially problematic) positive GIs, while even unveiling previously neglected negative GIs.
Figure 3: Evidence that the NetWF reduces experimental noise in the budding yeast S. cerevisiae genetic interaction (GI) network. (A) Validation of negative GI values against three biological benchmarks: protein–protein interaction (PPI), protein co-complex membership, and Gene Ontology (GO) bioprocess term co-annotation. Relative to the raw data, NetWF achieves higher area under the precision–recall curve (AUPRC) and larger fold enrichment across all benchmarks, while the optimal shrinker (OS) baseline delivers inconsistent performance. (B) Ten-fold cross validation (CV) test: in each fold, a random subset of GI entries is masked and predicted from the remaining data. The NetWF and the OS achieve significantly lower average mean squared error (MSE) than naïve mean imputation (MI), where the missing entries are imputed using the mean of available GI values in each gene. Error bars indicate standard errors across folds. (C) Example of inferring a novel GI within the exocyst complex (SEC15–SEC6) by the NetWF. Dashed edges highlight a few example profile similarities within the complex that facilitate information propagation, collectively shifting the GI value from zero to a strongly negative estimate. (D) Profile similarity networks (PSNs) showcasing similarities between gene pairs (Eq. 6), before and after the NetWF is applied to the raw GI network. Edge width scales with similarity, and similarities below $0.2$ are omitted for clarity.
More importantly, the filtered ExE network demonstrates stronger adherence to all three benchmarks (PPI, co-complex, GO bioprocess term co-annotation), improving both area under the precision–recall curve (AUPRC) and fold enrichment relative to the raw data (Figure 3A). Here, we compute AUPRC by ranking gene pairs from the most negative GI upward and synthesizing how effectively benchmark pairs are recovered as the selection threshold is progressively relaxed. Fold enrichment measures how overrepresented benchmark pairs are among the 1,0001{,}000 most negative GIs compared with their background prevalence in the overall ExE space.
@@
Next, to examine link prediction performance, we carry out a ten-fold cross validation test (Figure 3B). We divide the observed symmetric GI entries into 10 partitions, and in each fold, we mask the partitioned entries to missing values and infer them from the remaining observations; see Materials and Methods for details. The NetWF delivers substantially lower MSE than a naïve mean imputation that fills missing entries using the mean of available GI values for each gene.
@@
As a case study, we successfully recover a missing interaction from the input data within the exocyst complex [42]. The NetWF infers a strong negative GI between genes SEC15–SEC6 ($\epsilon=-0.21$), in agreement with the known biophysical interaction [23, 54]. Although the prediction originates from the NetWF processing collective information from the entire network, we can provide an interpretation within the exocyst subgraph (Figure 3C). Notably, SEC15 is highly similar to SEC10 and SEC8 in GI profile, while SEC6 is similar to SEC5; therefore, the strong negative GI of SEC10–SEC5 and SEC8–SEC5 contribute to the negative SEC15–SEC6 estimate. See SI Figure 3 for the physical structure of the exocyst complex, noting the SEC15–SEC6 physical contact.
@@
Finally, to quantify gene-level similarity, we construct PSNs in which edge weights are Pearson correlations between GI profiles (Eq. 6). Using the originally proposed cutoff >0.2>0.2 [19], the PSN has 7,5637{,}563 edges before filtering and 22,68522{,}685 edges after filtering (Figure 3D). This three-fold increase indicates fewer idiosyncratic differences in gene connection profiles, i.e. a smaller effective degrees of freedom. Together with the evidence that the biological signal is preserved or even enhanced (Figure 3A), this suggests a higher signal-to-noise ratio after filtering.
@@
In comparison, the OS baseline behaves inconsistently across evaluations. It adheres worse to the protein co-complex membership benchmarks than the raw data (Figure 3A), yet performs well in cross validation (Figure 3B), yields a dense PSN (58,63058{,}630 edges), and also recovers the SEC15–SEC6 interaction ($\epsilon=-0.15$).
@@
To test how much the performance of the NetWF depends on having access to the heterogeneous noise information, we run a restricted, homogenized-noise version in which we replace the heterogeneous noise variances by their mean, just like for the OS (SI Figure 4). The restricted results across the board change only marginally compared to those with the full NetWF, with over a $98\%$ overlap among the negative GIs. This result indicates that the NetWF can be a useful tool across various applications, even without detailed noise characterization, a point we further investigate in our next application.
@@
The Enron Corpus is a publicly available database containing internal email exchanges among key Enron employees in the years leading up to the company’s famous bankruptcy in December 2001, and it has become a widely used resource for studying organizational communication patterns [31]. We use this dataset for two complementary goals. First, we demonstrate the performance of the NetWF in the case of heterogeneous and correlated noise covariance $C_{n}$, the highest level of noise characteristics information. Second, we take the opportunity to embrace the more common empirical setting where no noise statistics is provided, and stress test the extent to which the NetWF can perform denoising under this far-from-optimal condition.
Figure 4: Evidence that the NetWF reduces sampling noise in the Enron Corpus email network. (A) Full-year email exchange frequency network of 2001 between employees as nodes. Directed edges point from sender to recipient, with edge weight equal to email frequency (defined as the number of emails per month). Node size reflects the sum of incoming and outgoing weights. For clarity, edges with weight <0.2<0.2 are not shown. (B) Raw October 2001 snapshot, treated as a noisy calendar-month sample of the full-year network, exhibiting pronounced fluctuations in edge weights. (C) October 2001 after applying the NetWF, with the extreme weights reduced without compromising the overall structure. (D) Relative to the raw October network and the OS baseline, the NetWF largely reduces the October MSE with respect to the full-year network. Repeating for all twelve calendar-month snapshots, the NetWF also attains the lowest and most stable average MSE. Error bars indicate standard errors across calendar months.